Server Status ­ Wave Source Design

. . . .
Server Status

Server Move

The WSD server is being moved into a new cabinet September 16th between 1AM and 8PM Eastern Time Zone (AEST). Total expected downtime will be 30 minutes. Thank you for your kind patience while we perform this maintenance. Connectivity for your websites will be greatly improved in the new cabinet and should provide overall better performance.

 

Server back online

All functions have now been restored, all websites are up and email functions have been restored. Once again, we apologise for any disruption to website services, and thank you for your patience.

Kernel update has failed

This is similar to an operating system upgrade. An automatic upgrade process for CentOS has been initiated for security reasons. However, some external dependencies have prevented the server from rebooting.

This unfortunately affects all services on the server - HTTP, email, MySQL etc.

We are rolling back the kernel to previous version which will take a little while, as the operating system files need to be rewritten with previous ones in backup...

DOS (Denial of Service) Attack

A DOS attack has been ongoing for the past 24 hours on a target server in our network cluster. While this does not affect sites on our server directly, network access to sites on our server is very slow leading to timeouts and non-connection messages in browsers.

The DOS attack has been closed down, but network conditions may persist for the next 12 hours, resolving to normal access by midday Monday...

Hacker exploit removed

Unexpected heavy load on server has been discovered to have been caused by a hacker exploiting a vulnerability on one user account on the server. User account has been temporarily suspended while we secure the website in question. Server loads are now down to their normal level, website access is back to normal for all client accounts.

Heavy load under investigation

We are encountering heavy loads on the system. Investigation has commenced to determine cause and solve the problem.

Normal traffic conditions now resuming

The network error has been resolved and normal traffic conditions are returning. Expect maybe a 5 to 10 minute transition period where access may be patchy as the datacentre catches up with the mail queue and page requests.

Traffic choking at node

We're currently seeing a condition where network traffic is being choked at a node at the data centre. Web pages are very slow to load, email connections will timeout. Investigation as to cuase is under way and we'll update as the situation changes.

Server migration complete

All hosting accounts have now completed their migration to our new server, WSD2, and are operating normally.